please help with ALL questions thank you1. Using excel, plot your standard curve based on Table 9-1 and calculate the line of best fit. Put concentration in the x coTable 8-1 Dilutions and corresponding concentration and OD values for yeast suspension #1 OD 600 YPD Broth (diluent) Yeast su
Answer #1

Load the data in Excel and insert a scatter plot, add a linear trendline, and then add equation using excel. The equation obtained in this case and graph is provided below.

Calibration curve Yeast y = 9E-08x +0.5237 1.2 1 0.8 0.D 600 nm 0.6 0.4 0.2 0 0 2000000 6000000 8000000 4000000 Concentration

Answers

1. The linear equation is in the form of y=mx+b

Slope (m)=9E-08

y-intercept (b)=0.5237

2. No, this standard can not be used for E. coli cells. A separate standard curve is required for E. coli bacteria. Since the growth curve of different bacteria and fungus is different we need separate curves to obtain results.

3. In order to calculate the unknown values, the y=mx+b can be used to determine the diluted concentration.

y=mx+b

x=\frac{y-b}{m}

x=\frac{y-0.5237}{9E-08}

4. Viable cells are the living cells in cell culture since only the living cells are able to form colonies on plates. Thus when we grow cells on the plate by spreading only live/viable cells will form colonies.

5. I think it is easy to carry out the turbidimetry method(O.D 600nm) since it requires less time and effort. The plate count method is a long term method.

6. The plate count method is more accurate in determining the original concentration since it provides only the viable cells. A seen from the data an O.D of 0.106 is obtained even when no viable cfu/mL is in the dilution. This method is able to differentiate between the live and dead cells and provide an accurate determination of the growth. While the turbidimetry method provides data for both dead and live cells.